# Large Language Models Understand Layout

### v1

@@ Abstract
Large language models (LLMs) demonstrate extraordinary abilities in a wide range of natural language processing (NLP) tasks. In this paper, we show that, beyond text understanding capability, LLMs are capable of processing text layouts that are denoted by spatial markers. They are able to answer questions that require explicit spatial perceiving and reasoning, while a drastic performance drop is observed when the spatial markers from the original data are excluded. We perform a series of experiments with the GPT-3.5, Baichuan2, Llama2 and ChatGLM3 models on various types of layout-sensitive datasets for further analysis. The experimental results reveal that the layout understanding ability of LLMs is mainly introduced by the coding data for pretraining, which is further enhanced at the instruction-tuning stage. In addition, layout understanding can be enhanced by integrating low-cost, auto-generated data approached by a novel text game. Finally, we show that layout understanding ability is beneficial for building efficient visual question-answering (VQA) systems.
@@ 1 Introduction
Beyond text understanding capability, we find that LLMs are capable of processing text layouts that are denoted by spatial markers. As shown in Figure 1, we conceptualize the newline-separated plain text as a "visual" two-dimensional canvas, as text editors and browsers are visually two-dimensional intuitively. Three identical questions with distinct answers are arranged in different orientations, interspersed with space markers (denoted as layout). We inquire with ChatGPT about the answers in various orientations. Remarkably, ChatGPT provides accurate responses, and some other open-source LLMs also demonstrate reasonable results. To compare with this, we exclude the space markers from the original data (denoted as strip), resulting in a substantial decline in performance.
@@
First, we build up a dataset called TextLayoutQA to evaluate LLMs’ text layout understanding capability. Through experiments with the GPT-3.5, Baichuan2 [45], Llama2 [39] and ChatGLM3 [48] models, we uncover that the incorporation of text layout information substantially enhances model performance, resulting in an 8$\sim$25gain compared to text without layout.
@@ 2 Related Work
In recent years, LLMs, represented by the GPT series, have demonstrated strong text comprehension abilities. In this domain, we are aware of some research efforts dedicated to exploring the performance of LLMs in spatial reasoning [47, 4], as well as the application of LLMs in graph analysis, understanding, and visualization [2, 5, 9, 35, 41, 50, 51].
@@
VQA is a task that involves answering questions about images, spanning various formats such as receipts, web pages, tables, documents, or even natural images containing textual information. This task fundamentally requires models to understand multiple modalities. Previous work predominantly focuses on multimodal pre-trained models [18, 12, 38, 11, 21], aiming to leverage all modal information. However, Liu et al. [25] point out that LLMs struggle with text-rich VQA scenarios. The main reasons include the short token lengths, usually 256, of the textual input from the visual encoder to the text encoder in multimodal LLMs, resulting in a significant loss of textual information. Additionally, the low resolution of the image encoder, typically 224*224, compresses and loses a considerable amount of textual information in text-rich images. Under the circumstances, recent work starts exploring the performance of purely textual LLMs in answering questions using only serialized text from images [10], leveraging the high accuracy of optical character recognition (OCR) models in recognizing long text.
@@ 3 Layout Understanding Based on Spatial Markers
Subsequently, we investigate whether the text layout understanding capability emerges during the pre-training stage or the instruction-tuning stage. Our hypothesis posits that the training corpora with consecutive spaces, including programming code, tables, HTML, YAML-formatted data, may contribute to the text layout understanding capability of LLMs. To validate this hypothesis, we construct an instruction-tuning dataset that does not include data with consecutive spaces as the instruction-basic dataset. We perform instruction-tuning on LLMs of different types and sizes, utilizing the characteristic of catastrophic forgetting to induce LLMs to "forget" text layout understanding capability. Subsequently, we add training corpora containing consecutive spaces, such as code and table corpora, to the instruction-basic dataset to observe whether the text layout understanding capability is recovered.
Furthermore, inspired by the well-known text game, word search puzzle, we devise a novel text game aiming at enhancing text layout comprehension through gameplay. Starting pre-training from scratch with different training corpora would be the most direct validation method. However, due to the substantial computational resources required, it is beyond the scope of our team’s capacity and could be considered as future work.
@@ 4.1 Datasets
In this section, we describe all the datasets we use in the experiments. These include three public datasets: XfundQA, FetaQA [27], DocVQA [26], along with a generated dataset for text layout understanding evaluation, named TextLayoutQA. Additionally, we propose various instruction-tuning datasets including instruction-basic, instruction-code, instruction-table, instruction-generated and instruction-test.
@@
A layout QA dataset generated specifically for testing the layout understanding capability of LLMs. This dataset revolves around enumerating items from various shopping lists arranged in random orientations within the text canvas. As illustrated in Figure 2a, some shopping lists are randomly positioned in four orientations (top-left, top-right, bottom-left, bottom-right) on a newline-separated plain text canvas filled with space markers. Each shopping list is assigned a name (A, B, C, or D) and comprises different products. Both the name and items within the same shopping list are first-letter aligned. For comparison, a version without layout information is constructed for each sample, which involves replacing consecutive space and newline markers with a single space marker. A minimum of two consecutive space markers are maintained between any two shopping lists. Figure 2b illustrates the "without layout" version corresponding to 2a. The paired samples, with and without layout, share the same set of three questions as shown in Figure 2c.
@@ A B footwear lenses movies walkers jet
GPTeacher 431 0.431
alpaca 414 0.414
webGPT 173 0.173
dolly 128 0.128
Auto-CoT 59 0.059
GAOKAO 17 0.017
@@
A testing dataset designed to assess the efficacy of instruction-tuning, consisting of four segments of data, as illustrated below, with a total of 800 samples. The evaluation metrics adopted for all instruction-tuning datasets are ROUGE-L and recall.
@@ 4.2 LLMs
In our experiments, we select several LLMs for evaluating TextLayoutQA, including GLM3, Llama2, Baichuan2, and GPT-3.5-Turbo. These models cover various sizes and types of open-source and proprietary LLMs that are currently popular.
@@
  • GPT-3.5-Turbo: launched by OpenAI, GPT-3.5-Turbo is an advanced language model derived from GPT-3, which has 175B parameters.
@@ 4.3.1 General Evaluation
Table 3 shows the evaluation results of different LLMs on TextLayoutQA with (layout) and without (strip) text layout information. Compared to the strip, various LLMs achieve a performance improvement of 8$\sim$25with text layout information, indicating the models’ ability to understand text alignment, layout, and orientation.
@@
ChatGLM3, Llama2, Baichuan2, and GPT-3.5-Turbo all use byte pair encoding (BPE) [34] from SentencePiece [20] as their tokenizer. We find that the tokenizers of these LLMs encode different lengths of consecutive spaces with distinct tokens. Table 4 illustrates the maximum lengths of consecutive punctuation that can be encoded by different LLMs’ tokenizers. GPT-3.5-Turbo has different tokens for most consecutive punctuation, while ChatGLM3, Llama2, and Baichuan2 only have different tokens for relatively limited consecutive punctuation. A commonality among them is that they all support encoding relatively long consecutive spaces, indicating that consecutive spaces have a certain proportion in the training corpora. Given that some programming languages like Python and certain file formats like YAML are sensitive to indentation, we assume that these corpora aid LLMs in grasping how consecutive spaces align texts, thereby acquiring text layout understanding capability.
@@
LLMs Space Tab Newline Exclamation Comma Full-stop ChatGLM3 15 1 1 2 1 4 Llama2 15 1 1 2 1 4 Baichuan2 32 1 1 1 2 3 GPT-3.5-Turbo 81 20 14 5 4 9
In TextLayoutQA, we use space and newline as spatial markers. For the ablation study, we investigate the other three characters as spatial markers: tab, caron (an accent mark), and a random vanilla character a. Notably, newline is still used to separate text lines. Table 5 delineates the effects of various LLMs on the TextLayoutQA dataset when deploying these characters as spatial markers. Generally, spaces consistently result in optimal performance for the majority of LLMs. The character a generally exhibits the poorest performance across various LLMs due to its lack of spatial semantics. In addition, LLMs do not generalize to interpret consecutive a as a spatial marker. Notably, although caron marker is rare in corpora, it still outperforms strip for most LLMs.
@@
4.3.2 The Origin of Layout Understanding Capability
Due to the lack of ability to strictly follow instructions, the outputs of the base models are difficult to align with the reference of the QA tasks. We therefore employ perplexity as the metric to ensure fair comparison of layout understanding capability between base and chat models. Perplexity is a widely used metric for assessing language models. Lower perplexity indicates better modeling performance. By comparing the perplexity of different LLMs on the TextLayoutQA dataset with and without text layout, the results presented in Table 6 are obtained. Notably, all the base models exhibit a lower perplexity on text with layout compared to text without layout, suggesting that the base models inherently possess some level of text layout understanding during the pre-training stage. Following instruction-tuning, all the chat models demonstrate a lower perplexity compared to the base models when processing text with layout as opposed to text without layout. This indicates that instruction-tuning further enhances the text layout understanding capability. It should be noted that, to mitigate the influence of context length on perplexity, newline markers are used for padding at the beginning of text without layout, with the padding length being the difference between the length of text without layout and the length of text with layout after tokenization.
@@ 4.3.2 The Origin of Layout Understanding Capability
Table 7 presents the training corpora utilized by various LLMs during the pre-training stage. Notably, the training corpora for GLM3 and Llama2 are not explicitly published, so related information about GLM-130b and Llama is considered. We find that GLM, Llama, and GPT-3 all use datasets such as CommonCrawl, Wikipedia, and Books (Pile includes CommonCrawl, Wikipedia, and Books) in their pre-training. CommonCrawl is a large-scale, unstructured, multilingual web dataset containing over 8 years of web crawler data. Additionally, GLM and Llama utilize code-related sources like GitHub and StackExchange. We do find some examples with various text layouts sourced from GitHub and StackExchange within the Pile dataset. The specific examples can be referred to Appendix.
@@
LLMs Training corpora GLM Pile, Chinese WudaoCorpora, Chinese corpora (including online forums, encyclopedia, and QA) crawled from the website Llama CommonCrawl [36], C4 [33], Github, Wikipedia, Books, ArXiv, StackExchange Baichuan2 General internet webpages, books, research papers, codebases, and more GPT-3 CommonCrawl, WebText [32], Books1, Books2, Wikipedia
We perform instruction-tuning on the instruction-basic, instruction-code, instruction-table and instruction-generated dataset using Firefly [46] tuning framework. Each dataset is partitioned into training and validation sets with a ratio of 98:2. The training sets undergo 5 epochs with early stopping. As expected, adding different instruction types improves the corresponding ability. For further information, please refer to Appendix.
Table 8 presents the performance of LLMs on TextLayoutQA following various types of instruction-tuning. Due to the characteristic of catastrophic forgetting, the text layout capability decreases after instruction-basic tuning compared to the chat model, except for the Llama2 series. This is because the Llama2 chat models tend to produce long responses and sometimes fail to follow the required output format. However, a significant recovery in text layout capability is observed after instruction-code tuning, underscoring the crucial role of code-related data in enhancing text layout understanding. In contrast, the model fine-tuned on instruction-table experiences a decrease in text layout capability, indicating that the table-related data do not contribute to the text layout capability. It is noteworthy that performance shows enhancement following instruction-generated tuning. For specific small models like Llama2-7B and Baichuan2-7B, instruction-generated tuning even yields the best result among all instruction-tuning datasets. Considering that generating data is significantly more convenient and cost-effective than collecting, we have laid a promising path for pre-training LLMs.
@@ 4.3.3 Applications
In this section, we illustrate the utilization of LLMs’ text layout understanding capability in the text-rich VQA domain. We introduce a method named textLayoutParser designed to parse texts with diverse layouts from documents, including plain texts, forms, tables, images, and their combinations. The method involves the placement of text on a two-dimensional character canvas according to the text’s coordinates. Detailed implementation is available in Appendix. We evaluate the zero-shot performance on the test sets of three datasets—XfundQA, DocVQA, and FeTaQA. The prompts utilized for each dataset are provided in Appendix.
@@
We use the OCR output provided by the dataset and construct corpora with text layout using textLayoutParser. As a comparison, we replace consecutive spaces and newlines with a single space marker, forming corpora without text layout. The evaluation results of different LLMs on XfundQA with and without text layout are presented in Table 11. Notably, corpora with text layout lead to performance improvements ranging from 1.96% to 9.55% compared to corpora without text layout.
@@
We use the OCR output provided by the dataset and construct corpora with text layout using textLayoutParser. For comparison, consecutive spaces and newlines are replaced with a single space marker, forming corpora without text layout. Table 12 shows the evaluation results of different LLMs on the DocVQA test set with and without text layout. Compared to corpora without text layout, different LLMs achieved performance improvements of 2.67% to 4.27% on corpora with text layout.
@@
Difference text layout encoding methods are tailored to specific cases. For instance, in the context of table QA, common table encoding techniques include employing identifiers to distinguish headers and rows (referred to as Linear) [23, 15] and representing each element as a col-row-value triplet to create a list (referred to as Triplet) [40]. Apart from our proposed method, we explore several other text layout encoding techniques for an ablation study. Examples of different table encoding methods can be referred to Appendix. Table 4.3.3 provides a performance assessment of various table encoding methods on the FeTaQA test set. Our proposed method outperforms others for ChatGLM3-6B, Llama2-7B, and GPT-3.5-Turbo. Conversely, for Baichuan2-13B, the Linear encoding method demonstrates superior results.

### v3

@@ Abstract
Large language models (LLMs) demonstrate extraordinary abilities in a wide range of natural language processing (NLP) tasks. In this paper, we show that, beyond text understanding capability, LLMs are capable of processing text layouts that are denoted by spatial markers. They are able to answer questions that require explicit spatial perceiving and reasoning, while a drastic performance drop is observed when the spatial markers from the original data are excluded. We perform a series of experiments with the GPT-3.5/4, Baichuan2, Llama2, and ChatGLM3 models on various types of layout-sensitive datasets for further analysis. The experimental results reveal that the layout understanding ability of LLMs is mainly introduced by the coding data for pre-training, which is further enhanced at the instruction-tuning stage. In addition, layout understanding can be enhanced by integrating low-cost, auto-generated data approached by a novel text game. Finally, we show that layout understanding ability is beneficial for building efficient visual question-answering (VQA) systems.
@@ 1 Introduction
Beyond text understanding capability, we find that LLMs are capable of processing text layouts that are denoted by spatial markers. As shown in Figure 1, we conceptualize the newline-separated plain text as a "visual" two-dimensional canvas, as text editors and browsers are visually two-dimensional intuitively. Three identical questions with distinct answers are arranged in different orientations, interspersed with space markers (denoted as layout). We inquire with ChatGPT about the answers to various orientations. Remarkably, ChatGPT provides accurate responses, and some other open-source LLMs also demonstrate reasonable results. To compare with this, we exclude the space markers from the original data (denoted as strip), resulting in a substantial decline in performance. More examples can be found in Appendix A.
@@
First, we build up a dataset called TextLayoutQA to evaluate LLMs’ text layout understanding capability. Through experiments with the GPT-3.5/4, Baichuan2 [45], Llama2 [39] and ChatGLM3 [48] models, we uncover that the incorporation of text layout information substantially enhances model performance, resulting in an 8$\sim$33gain compared to text without layout.
@@ 2 Related Work
In recent years, LLMs, represented by the GPT series, have demonstrated strong text comprehension abilities. In this domain, we are aware of some research efforts dedicated to exploring the performance of LLMs in spatial reasoning [47, 4], as well as the application of LLMs in graph analysis, understanding, and visualization [2, 5, 9, 35, 41].
@@
VQA is a task that involves answering questions about images, spanning various formats such as receipts, web pages, tables, documents, or even natural images containing textual information. This task fundamentally requires models to understand multiple modalities. Previous work predominantly focuses on multimodal pre-trained models [18, 12, 38, 11, 21], aiming to leverage all modal information. However, Liu et al. [25] point out that LLMs struggle with text-rich VQA scenarios. The main reasons include the short token lengths, usually 256, of the textual input from the visual encoder to the text encoder in multimodal LLMs, resulting in a significant loss of textual information. Additionally, the low resolution of the image encoder, typically 224*224, compresses and loses a considerable amount of textual information in text-rich images. Under the circumstances, recent work has started exploring the performance of purely textual LLMs in answering questions using only serialized text from images [10], leveraging the high accuracy of optical character recognition (OCR) models in recognizing long text.
@@ 3 Layout Understanding Based on Spatial Markers
Subsequently, we investigate whether the text layout understanding capability emerges during the pre-training stage or the instruction-tuning stage. Our hypothesis posits that the training corpora with consecutive spaces, including programming code, tables, HTML, and YAML-formatted data, may contribute to the text layout understanding capability of LLMs. To validate this hypothesis, we construct an instruction-tuning dataset that does not include data with consecutive spaces as the instruction-basic dataset. We perform instruction-tuning on LLMs of different types and sizes, utilizing the characteristic of catastrophic forgetting to induce LLMs to "forget" text layout understanding capability. Subsequently, we add training corpora containing consecutive spaces, such as code and table corpora, to the instruction-basic dataset to observe whether the text layout understanding capability is recovered.
Furthermore, inspired by the well-known text game, word search puzzle, we devise a novel text game aiming at enhancing text layout comprehension through gameplay. Starting pre-training from scratch with different training corpora would be the most direct validation method. However, due to the substantial computational resources required, it is beyond the scope of our team’s capacity and could be considered future work.
@@ 4.1 Datasets
In this section, we describe all the datasets we use in the experiments. These include three public datasets: XfundQA, FetaQA [27], DocVQA [26], along with a generated dataset for text layout understanding evaluation, named TextLayoutQA. Additionally, we propose various instruction-tuning datasets including instruction-basic, instruction-code, instruction-table, instruction-generated, and instruction-test.
@@
A layout QA dataset we proposed to test the layout understanding capability of LLMs. This dataset revolves around enumerating items from various shopping lists arranged in random orientations within the text canvas. As illustrated in Figure 2a, some shopping lists are randomly positioned in four orientations (top-left, top-right, bottom-left, bottom-right) on a newline-separated plain text canvas filled with space markers. Each shopping list is assigned a name (A, B, C, or D) and comprises different products. Both the name and items within the same shopping list are first-letter aligned. For comparison, a version without layout information is constructed for each sample, which involves replacing consecutive space and newline markers with a single space marker. A minimum of two consecutive space markers are maintained between any two shopping lists. Figure 2b illustrates the "without layout" version corresponding to 2a. The paired samples, with and without layout, share the same set of three questions as shown in Figure 2c.
@@ A B footwear lenses movies walkers jet
GPTeacher 431 0.43
alpaca 414 0.41
webGPT 173 0.17
dolly 128 0.12
Auto-CoT 59 0.05
GAOKAO 17 0.01
@@
A testing dataset designed to assess the efficacy of instruction tuning, consisting of four segments of data, as illustrated below, with a total of 800 samples. The evaluation metrics adopted for all instruction-tuning datasets are ROUGE-L and recall.
@@ 4.2 LLMs
In our experiments, we select several LLMs for evaluating TextLayoutQA, including GLM3, Llama2, Baichuan2, and GPT-3.5/4. These models cover various sizes and types of open-source and proprietary LLMs that are currently popular.
@@
  • GPT-3.5/4: developed by OpenAI, we utilize the GPT-3.5-Turbo and GPT-4-Turbo versions in our experiments. GPT-3.5 is a highly capable model known for generating human-like text based on the input provided. GPT-4 enhances this capability, delivering improved accuracy, comprehension, and contextual understanding.
@@ 4.3.1 General Evaluation
Table 3 shows the evaluation results of different LLMs on TextLayoutQA with (layout) and without (strip) text layout information. Compared to the strip, various LLMs achieve a performance improvement of 8$\sim$33with text layout information, indicating the models’ ability to understand text alignment, layout, and orientation.
@@
GPT-4-Turbo 65.42 98.83 +33.41
ChatGLM3, Llama2, Baichuan2, and GPT-3.5/4 all use byte pair encoding (BPE) [34] from SentencePiece [20] as their tokenizer. These tokenizers encode different lengths of consecutive spaces with distinct tokens. Table 4 shows that GPT-3.5/4 has distinct tokens for most consecutive punctuation, while ChatGLM3, Llama2, and Baichuan2 only do so for a limited range. However, all these models support encoding relatively long consecutive spaces, suggesting that such spaces are prevalent in their training data. Given that some programming languages like Python and certain file formats like YAML are sensitive to indentation, we assume that these corpora aid LLMs in grasping how consecutive spaces align texts, thereby acquiring text layout understanding capability.
@@
LLMs Space Tab Newline Exclamation Comma Full-stop ChatGLM3 15 1 1 2 1 4 Llama2 15 1 1 2 1 4 Baichuan2 32 1 1 1 2 3 GPT-3.5/4 81 20 14 5 4 9
In TextLayoutQA, we use space and newline as spatial markers. For the ablation study, we investigate another three characters as spatial markers: tab, caron (an accent mark), and a random vanilla character a. Notably, newline is still used to separate text lines. Table 5 delineates the effects of various LLMs on the TextLayoutQA dataset when deploying these characters as spatial markers. Generally, spaces consistently result in optimal performance for the majority of LLMs. The character a generally exhibits the poorest performance across various LLMs due to its lack of spatial semantics. Notably, although caron marker is rare in corpora, it still outperforms strip for most LLMs.
4.3.2 The Origin of Layout Understanding Capability
@@
GPT-4-Turbo 65.42 98.83 98.82 93.90 85.37
Due to the base models’ limited ability to strictly follow instructions, their outputs often misalign with QA task references. To ensure a fair comparison of layout understanding between base and chat models, we use perplexity, a common metric for assessing language models, where lower perplexity indicates better performance. Table 6 compares the perplexity of different LLMs on the TextLayoutQA dataset with and without layout. All base models show lower perplexity on text with layout, suggesting some inherent text layout understanding from pre-training. After instruction-tuning, chat models demonstrate even lower perplexity on text with layout compared to base models, indicating enhanced layout understanding. To minimize the influence of context length on perplexity, newline markers are used for padding at the beginning of text without layout, with the padding length being the difference between the length of text without layout and the length of text with layout after tokenization.
@@
Table 7 presents the training corpora utilized by various LLMs during the pre-training stage. Notably, the training corpora for GLM3, Llama2, and GPT-4 are not explicitly published, so related information about GLM-130b, Llama, and GPT-3.5 is considered respectively. We find that GLM, Llama, and GPT-3.5 all use datasets such as CommonCrawl, Wikipedia, and Books (Pile includes CommonCrawl, Wikipedia, and Books) in their pre-training. CommonCrawl is a large-scale, unstructured, multilingual web dataset containing over 8 years of web crawler data. Additionally, GLM and Llama utilize code-related sources like GitHub and StackExchange. We do find some examples with various text layouts sourced from GitHub and StackExchange within the Pile dataset. The specific examples can be found in Appendix B.
@@
LLMs Training corpora GLM Pile, Chinese WudaoCorpora, Chinese corpora (including online forums, encyclopedia, and QA) crawled from the website Llama CommonCrawl [36], C4 [33], Github, Wikipedia, Books, ArXiv, StackExchange Baichuan2 General internet webpages, books, research papers, codebases, and more GPT-3.5 CommonCrawl, WebText [32], Books1, Books2, Wikipedia
We perform instruction-tuning on the instruction-basic, instruction-code, instruction-table, and instruction-generated dataset using Firefly [46] tuning framework. Each dataset is partitioned into training and validation sets with a ratio of 98:2. The training sets undergo 5 epochs with early stopping. As expected, adding different instruction types improves the corresponding ability. The instruction-tuning parameters can be referred to Appendix C.
Table 8 presents the performance of LLMs on TextLayoutQA following various types of instruction-tuning. The comprehensive result can be found in Appendix D. Due to the characteristic of catastrophic forgetting, the text layout capability decreases after instruction-basic tuning compared to the chat model, except for the Llama2 series. This is because the Llama2 chat models tend to produce long responses and sometimes fail to follow the required output format. However, a significant recovery in text layout capability is observed after instruction-code tuning, underscoring the crucial role of code-related data in enhancing text layout understanding. In contrast, the model fine-tuned on instruction-table data experiences a decrease in text layout capability, indicating that the table-related data do not contribute to the text layout capability. It is noteworthy that performance shows enhancement following instruction-generated tuning. For specific small models like Llama2-7B and Baichuan2-7B, instruction-generated tuning even yields the best result among all instruction-tuning datasets. Considering that generating data is significantly more convenient and cost-effective than collecting, we have laid a promising path for pre-training LLMs.
@@ 4.3.3 Applications
In this section, we illustrate the utilization of LLMs’ text layout understanding capability in the text-rich VQA domain. We introduce a method named textLayoutParser designed to parse texts with diverse layouts from documents, including plain texts, forms, tables, images, and their combinations. The method involves the placement of text on a two-dimensional character canvas according to the text’s coordinates. Detailed implementation is available in Appendix E. We evaluate the zero-shot performance on the test sets of three datasets: XfundQA, DocVQA, and FeTaQA. The prompts utilized for each dataset are provided in Appendix F.
@@
We use the OCR output provided by the dataset and construct corpora with text layout using textLayoutParser. As a comparison, we replace consecutive spaces and newlines with a single space marker, forming corpora without text layout. The evaluation results of different LLMs on XfundQA with and without text layout are presented in Table 11. Notably, corpora with text layout lead to performance improvements ranging from 0.83% to 9.55% compared to corpora without text layout. This infers that the LLM’s text layout understanding capability extends to Chinese.
@@
We use the OCR output provided by the dataset and construct corpora with text layout using textLayoutParser. For comparison, consecutive spaces and newlines are replaced with a single space marker, forming corpora without text layout. Table 12 shows the evaluation results of different LLMs on the DocVQA test set with and without text layout. Compared to corpora without text layout, different LLMs achieved performance improvements of 2.67% to 4.63% on corpora with text layout.
@@
Difference text layout encoding methods are tailored to specific cases. For instance, in the context of table QA, common table encoding techniques include employing identifiers to distinguish headers and rows (referred to as Linear) [23, 15] and representing each element as a col-row-value triplet to create a list (referred to as Triplet) [40]. Apart from our proposed method, we explore several other text layout encoding techniques for an ablation study. Examples of different table encoding methods are available in Appendix G. Table 4.3.3 provides a performance assessment of various table encoding methods on the FeTaQA test set. Our proposed method outperforms others for ChatGLM3-6B, Llama2-7B, and GPT-3.5-Turbo. Conversely, for Baichuan2-13B and GPT-4-Turbo, the Linear encoding method demonstrates superior results.